# Comparative Study of Recurrent Neural Networks for Virtual Analog Audio Effects Modeling

### v3

@@ 4 Results
Table 3: MAE, ESR, $M_{RMSE}$, and $M_{SF}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
Dataset Model MAE ESR $M_{RMSE}$ $M_{SF}$
OD300 LSTM $\boldsymbol{1.14\cdot 10^{2}}$ $\boldsymbol{1.06\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $1.60\cdot 10^{-2}$
ED $1.40\cdot 10^{-2}$ $1.54\cdot 10^{-1}$ $1.30\cdot 10^{-2}$ $1.19\cdot 10^{-2}$
LRU $1.73\cdot 10^{-2}$ $2.35\cdot 10^{-1}$ $1.65\cdot 10^{-2}$ $1.77\cdot 10^{-2}$
S4D $1.29\cdot 10^{-2}$ $1.40\cdot 10^{-1}$ $\boldsymbol{1.20\cdot 10^{-2}}$ $\boldsymbol{8.50\cdot 10^{-3}}$
Neutron’s LSTM $\boldsymbol{2.02\cdot 10^{-2}}$ $\boldsymbol{2.50\cdot 10^{-1}}$ $\boldsymbol{1.86\cdot 10^{-2}}$ $5.74\cdot 10^{-3}$
OD ED $2.54\cdot 10^{-2}$ $3.07\cdot 10^{-1}$ $2.36\cdot 10^{-2}$ $6.33\cdot 10^{-3}$
LRU $2.14\cdot 10^{-2}$ $2.70\cdot 10^{-1}$ $1.94\cdot 10^{-2}$ $6.64\cdot 10^{-3}$
S4D $2.40\cdot 10^{-2}$ $2.88\cdot 10^{-1}$ $2.21\cdot 10^{-2}$ $\boldsymbol{5.71\cdot 10^{-3}}$
Helper LSTM $1.12\cdot 10^{-2}$ $2.67\cdot 10^{-1}$ $1.05\cdot 10^{-2}$ $2.78\cdot 10^{-2}$
Saturator ED $1.15\cdot 10^{-2}$ $2.58\cdot 10^{-1}$ $1.06\cdot 10^{-2}$ $1.29\cdot 10^{-2}$
LRU $\boldsymbol{1.04\cdot 10^{-2}}$ $3.46\cdot 10^{-1}$ $\boldsymbol{9.72\cdot 10^{-3}}$ $1.42\cdot 10^{-2}$
S4D $1.13\cdot 10^{-2}$ $\boldsymbol{2.57\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $\boldsymbol{1.22\cdot 10^{-2}}$
Pultec LSTM $1.36\cdot 10^{-2}$ $\boldsymbol{2.15\cdot 10^{-1}}$ $1.29\cdot 10^{-2}$ $8.85\cdot 10^{-3}$
Passive ED $1.40\cdot 10^{-2}$ $3.22\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $2.49\cdot 10^{-3}$
EQ LRU $\boldsymbol{1.32\cdot 10^{-2}}$ $3.00\cdot 10^{-1}$ $\boldsymbol{1.16\cdot 10^{-2}}$ $1.99\cdot 10^{-3}$
S4D $1.36\cdot 10^{-2}$ $3.08\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $\boldsymbol{1.79\cdot 10^{-3}}$
LA-2A LSTM $4.77\cdot 10^{-3}$ $1.03\cdot 10^{-1}$ $4.51\cdot 10^{-3}$ $1.12\cdot 10^{-2}$
ED $4.32\cdot 10^{-3}$ $8.91\cdot 10^{-2}$ $4.10\cdot 10^{-3}$ $\boldsymbol{7.28\cdot 10^{-3}}$
LRU $5.10\cdot 10^{-3}$ $1.06\cdot 10^{-1}$ $4.67\cdot 10^{-3}$ $7.84\cdot 10^{-3}$
S4D $\boldsymbol{3.70\cdot 10^{-3}}$ $\boldsymbol{7.87\cdot 10^{-2}}$ $\boldsymbol{3.51\cdot 10^{-3}}$ $1.00\cdot 10^{-2}$
CL 1B LSTM $1.06\cdot 10^{-2}$ $2.92\cdot 10^{-1}$ $9.33\cdot 10^{-3}$ $5.21\cdot 10^{-3}$
ED $\boldsymbol{4.25\cdot 10^{-3}}$ $\boldsymbol{1.46\cdot 10^{-1}}$ $\boldsymbol{4.05\cdot 10^{-3}}$ $\boldsymbol{4.18\cdot 10^{-3}}$
LRU $6.59\cdot 10^{-3}$ $1.88\cdot 10^{-2}$ $6.38\cdot 10^{-3}$ $5.49\cdot 10^{-3}$
S4D $5.19\cdot 10^{-3}$ $2.00\cdot 10^{-1}$ $5.02\cdot 10^{-3}$ $6.12\cdot 10^{-3}$
Neutron’s LSTM $3.90\cdot 10^{-2}$ $1.06$ $3.15\cdot 10^{-2}$ $3.03\cdot 10^{-2}$
Filter ED $2.51\cdot 10^{-2}$ $8.77\cdot 10^{-1}$ $2.18\cdot 10^{-2}$ $2.21\cdot 10^{-2}$
LRU $\boldsymbol{2.11\cdot 10^{-2}}$ $\boldsymbol{6.58\cdot 10^{-1}}$ $\boldsymbol{1.66\cdot 10^{-2}}$ $\boldsymbol{1.95\cdot 10^{-2}}$
S4D $2.22\cdot 10^{-2}$ $7.18\cdot 10^{-1}$ $1.97\cdot 10^{-2}$ $2.00\cdot 10^{-2}$
[…]
Table 4: $M_{STFT-1}$, $M_{STFT-2}$, and $M_{MFCC}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
Dataset Model $M_{STFT-1}$ $M_{STFT-2}$ $M_{MFCC}$
OD300 LSTM $4.53\cdot 10^{-1}$ $3.56\cdot 10^{-1}$ $6.08\cdot 10^{-1}$
ED $4.47\cdot 10^{-1}$ $3.52\cdot 10^{-1}$ $5.80\cdot 10^{-1}$
LRU $4.65\cdot 10^{-1}$ $4.55\cdot 10^{-1}$ $1.05$
S4D $\boldsymbol{4.19\cdot 10^{-1}}$ $\boldsymbol{3.11\cdot 10^{-1}}$ $\boldsymbol{5.38\cdot 10^{-1}}$
Neutron’s LSTM $5.79\cdot 10^{-1}$ $4.80\cdot 10^{-1}$ $3.49\cdot 10^{-1}$
OD ED $6.68\cdot 10^{-1}$ $5.64\cdot 10^{-1}$ $3.80\cdot 10^{-1}$
LRU $\boldsymbol{3.56\cdot 10^{-1}}$ $\boldsymbol{3.40\cdot 10^{-1}}$ $\boldsymbol{3.39\cdot 10^{-1}}$
S4D $6.40\cdot 10^{-1}$ $5.20\cdot 10^{-1}$ $3.59\cdot 10^{-1}$
Helper LSTM $6.56\cdot 10^{-1}$ $5.97\cdot 10^{-1}$ $4.72\cdot 10^{-1}$
Saturator ED $6.65\cdot 10^{-1}$ $\boldsymbol{5.33\cdot 10^{-1}}$ $4.14\cdot 10^{-1}$
LRU $\boldsymbol{5.62\cdot 10^{-1}}$ $5.44\cdot 10^{-1}$ $\boldsymbol{2.46\cdot 10^{-1}}$
S4D $6.73\cdot 10^{-1}$ $5.41\cdot 10^{-1}$ $3.67\cdot 10^{-1}$
Pultec LSTM $1.10$ $9.33\cdot 10^{-1}$ $2.06\cdot 10^{-1}$
Passive ED $6.74\cdot 10^{-1}$ $4.94\cdot 10^{-1}$ $1.91\cdot 10^{-1}$
EQ LRU $\boldsymbol{1.82\cdot 10^{-1}}$ $\boldsymbol{1.73\cdot 10^{-1}}$ $9.81\cdot 10^{-2}$
S4D $6.04\cdot 10^{-1}$ $4.09\cdot 10^{-1}$ $\boldsymbol{9.44\cdot 10^{-2}}$
LA-2A LSTM $4.68\cdot 10^{-1}$ $4.37\cdot 10^{-1}$ $3.10\cdot 10^{-1}$
ED $\boldsymbol{3.12\cdot 10^{-1}}$ $\boldsymbol{2.30\cdot 10^{-1}}$ $\boldsymbol{2.30\cdot 10^{-1}}$
LRU $2.75\cdot 10^{-1}$ $2.70\cdot 10^{-1}$ $2.93\cdot 10^{-1}$
S4D $3.19\cdot 10^{-1}$ $2.52\cdot 10^{-1}$ $2.53\cdot 10^{-1}$
CL 1B LSTM $4.70\cdot 10^{-1}$ $3.56\cdot 10^{-1}$ $2.26\cdot 10^{-1}$
ED $4.89\cdot 10^{-1}$ $4.17\cdot 10^{-1}$ $\boldsymbol{2.09\cdot 10^{-1}}$
LRU $4.39\cdot 10^{-1}$ $4.38\cdot 10^{-1}$ $4.52\cdot 10^{-1}$
S4D $\boldsymbol{2.95\cdot 10^{-1}}$ $\boldsymbol{2.79\cdot 10^{-1}}$ $2.93\cdot 10^{-1}$
Neutron’s LSTM $9.44\cdot 10^{-1}$ $9.33\cdot 10^{-1}$ $6.57\cdot 10^{-1}$
Filter ED $9.58\cdot 10^{-1}$ $9.38\cdot 10^{-1}$ $6.32\cdot 10^{-1}$
LRU $\boldsymbol{7.89\cdot 10^{-1}}$ $\boldsymbol{7.51\cdot 10^{-1}}$ $\boldsymbol{5.98\cdot 10^{-1}}$
S4D $9.13\cdot 10^{-1}$ $9.01\cdot 10^{-1}$ $6.28\cdot 10^{-1}$
[…]
Generally, performance metrics are relatively close across models across most effects examined in this study. Although the S4D models slightly outperform other models based on the metrics alone, it appears to be the most suitable architecture analog effect modeling, at least considering those within the scope of this study. The ED, LRU, and particularly the S4D models are more effective at encoding past signal information within their internal states. This is suggested by the case of the cases of optical compressors, where they significantly outperform other models. Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical time dependencies, such as overdrive, equalizer, and filters. Moreover, LSTM models tend to predict energy in the low-frequency spectrum more accurately.
@@ 5 Conclusion
The computational challenges of simulating the nonlinearities of analog circuits, described by complex differential equations, are a significant obstacle for real-time applications. The complexity of these nonlinearities makes them difficult to model accurately, but they also contribute to the unique sound characteristics of the devices, making them appealing for both academic and industrial purposes. Modeling these nonlinear aspects is a difficult task that requires simplifications and assumptions to ensure stability and efficiency in the numerical schemes. Artificial neural networks (ANN) have emerged as a popular data-driven approach for this modeling process. While they offer a convenient and automated approach, they may struggle with the computational complexity and flexibility required for handling control parameters. In some cases, hybrid solutions that incorporate physics knowledge or measurements from internal components may provide advantages. However, there are situations where black-box models are necessary and feasible. Recurrent neural networks (RNNs) utilize internal states as memory mechanisms, allowing them to encode the signal history. This feature enables more interactive solutions as their memory capacity is not solely dependent on the input amount. In this paper, we present a comparative analysis that examines the performance of the widely used Long Short-Term Memory (LSTM) models against recently proposed recurrent models, namely Linear Recurrent Unit (LRU) and State-Sapce models (SSM), which have demonstrated remarkable results in sequence modeling. To carry out this study, we design four models that incorporate parameter conditioning and emphasize low-latency response. The control parameters are integrated into the models using the Feature-wise Linear Modulation (FiLM) method. Our selection of effects includes overdrive, saturation, low pass filter, equalization, and compressors. We evaluate these models using various metrics, which provide valuable insights for future research and development. Specifically, our metrics consider the root-mean-squared energy envelopes, the transients, and the frequency content.
LSTM-based networks excel in replicating distortion and equalizer effects, showing their ability to encode the audio dependencies enough for these specific effects effectively. On the other hand, LSTM, when designed in an encoder-decoder fashion, LRUs and, in particular, SSMs, outperform other models when emulating saturation and compression. In the case of compression, they demonstrate their advantage in accurately capturing the signal’s history. Generally, SSMs have also proved to be an appropriate candidate technique in VA, although the gap with the other models is not always so pronounced.
@@
Conditioning techniques and performances should be further investigated in future work. The current conditioning blocks rely on affine transformation. Still, devices such as distortion or saturation could benefit from more complex transformations to account for the possible nonlinear influence of the parameters.

### v5

@@ 4 Results
Table 4: ESR, $M_{RMSE}$, $M_{SF}$, and $M_{STFT}$ metrics computed on the test set. The test sets include all conditioning values, if applicable.
Dataset Model ESR $M_{RMSE}$ $M_{SF}$ $M_{STFT}$ $M_{MFCC}$
OD300 LSTM $\boldsymbol{1.06\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $1.60\cdot 10^{-2}$ $3.56\cdot 10^{-1}$ $6.08\cdot 10^{-1}$
Overdrive ED $1.54\cdot 10^{-1}$ $1.30\cdot 10^{-2}$ $1.19\cdot 10^{-2}$ $3.52\cdot 10^{-1}$ $5.80\cdot 10^{-1}$
LRU $2.35\cdot 10^{-1}$ $1.65\cdot 10^{-2}$ $1.77\cdot 10^{-2}$ $4.55\cdot 10^{-1}$ $1.05$
S4D $1.40\cdot 10^{-1}$ $\boldsymbol{1.20\cdot 10^{-2}}$ $\boldsymbol{8.50\cdot 10^{-3}}$ $3.11\cdot 10^{-1}$ $\boldsymbol{5.38\cdot 10^{-1}}$
S6 $1.46\cdot 10^{-1}$ $1.25\cdot 10^{-2}$ $1.04\cdot 10^{-2}$ $\boldsymbol{2.78\cdot 10^{-1}}$ $5.61\cdot 10^{-1}$
Neutron LSTM $\boldsymbol{2.50\cdot 10^{-1}}$ $\boldsymbol{1.86\cdot 10^{-2}}$ $5.74\cdot 10^{-3}$ $4.80\cdot 10^{-1}$ $3.49\cdot 10^{-1}$
Overdrive ED $3.07\cdot 10^{-1}$ $2.36\cdot 10^{-2}$ $6.33\cdot 10^{-3}$ $5.64\cdot 10^{-1}$ $3.80\cdot 10^{-1}$
LRU $2.70\cdot 10^{-1}$ $1.94\cdot 10^{-2}$ $6.64\cdot 10^{-3}$ $\boldsymbol{3.40\cdot 10^{-1}}$ $3.39\cdot 10^{-1}$
S4D $2.88\cdot 10^{-1}$ $2.21\cdot 10^{-2}$ $5.71\cdot 10^{-3}$ $5.20\cdot 10^{-1}$ $3.59\cdot 10^{-1}$
S6 $2.65\cdot 10^{-1}$ $1.89\cdot 10^{-2}$ $\boldsymbol{4.97\cdot 10^{-3}}$ $3.86\cdot 10^{-1}$ $\boldsymbol{3.05\cdot 10^{-1}}$
Helper LSTM $2.67\cdot 10^{-1}$ $1.05\cdot 10^{-2}$ $2.78\cdot 10^{-2}$ $5.97\cdot 10^{-1}$ $4.72\cdot 10^{-1}$
Saturator ED $2.58\cdot 10^{-1}$ $1.06\cdot 10^{-2}$ $1.29\cdot 10^{-2}$ $5.33\cdot 10^{-1}$ $4.14\cdot 10^{-1}$
LRU $3.46\cdot 10^{-1}$ $\boldsymbol{9.72\cdot 10^{-3}}$ $1.42\cdot 10^{-2}$ $5.44\cdot 10^{-1}$ $\boldsymbol{2.46\cdot 10^{-1}}$
S4D $\boldsymbol{2.57\cdot 10^{-1}}$ $1.06\cdot 10^{-2}$ $\boldsymbol{1.22\cdot 10^{-2}}$ $5.41\cdot 10^{-1}$ $3.67\cdot 10^{-1}$
S6 $2.65\cdot 10^{-1}$ $1.01\cdot 10^{-2}$ $1.31\cdot 10^{-2}$ $\boldsymbol{4.28\cdot 10^{-1}}$ $3.88\cdot 10^{-1}$
Pultec LSTM $\boldsymbol{2.15\cdot 10^{-1}}$ $1.29\cdot 10^{-2}$ $8.85\cdot 10^{-3}$ $9.33\cdot 10^{-1}$ $2.06\cdot 10^{-1}$
Passive ED $3.22\cdot 10^{-1}$ $1.24\cdot 10^{-2}$ $2.49\cdot 10^{-3}$ $4.94\cdot 10^{-1}$ $1.91\cdot 10^{-1}$
Equalizer LRU $3.00\cdot 10^{-1}$ $\boldsymbol{1.16\cdot 10^{-2}}$ $1.99\cdot 10^{-3}$ $\boldsymbol{1.73\cdot 10^{-1}}$ $9.81\cdot 10^{-2}$
S4D $3.08\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $\boldsymbol{1.79\cdot 10^{-3}}$ $4.09\cdot 10^{-1}$ $\boldsymbol{9.44\cdot 10^{-2}}$
S6 $2.80\cdot 10^{-1}$ $1.19\cdot 10^{-2}$ $3.89\cdot 10^{-3}$ $2.87\cdot 10^{-1}$ $1.26\cdot 10^{-1}$
LA-2A LSTM $1.03\cdot 10^{-1}$ $4.51\cdot 10^{-3}$ $1.12\cdot 10^{-2}$ $4.37\cdot 10^{-1}$ $3.10\cdot 10^{-1}$
Optical ED $8.91\cdot 10^{-2}$ $4.10\cdot 10^{-3}$ $7.28\cdot 10^{-3}$ $2.30\cdot 10^{-1}$ $2.30\cdot 10^{-1}$
Compressor LRU $1.06\cdot 10^{-1}$ $4.67\cdot 10^{-3}$ $7.84\cdot 10^{-3}$ $2.70\cdot 10^{-1}$ $2.93\cdot 10^{-1}$
S4D $7.87\cdot 10^{-2}$ $\boldsymbol{3.51\cdot 10^{-3}}$ $1.00\cdot 10^{-2}$ $2.52\cdot 10^{-1}$ $2.53\cdot 10^{-1}$
S6 $\boldsymbol{7.71\cdot 10^{-2}}$ $3.63\cdot 10^{-3}$ $\boldsymbol{6.26\cdot 10^{-3}}$ $\boldsymbol{1.84\cdot 10^{-1}}$ $\boldsymbol{2.12\cdot 10^{-1}}$
CL 1B LSTM $2.92\cdot 10^{-1}$ $9.33\cdot 10^{-3}$ $5.21\cdot 10^{-3}$ $3.56\cdot 10^{-1}$ $2.26\cdot 10^{-1}$
Optical ED $\boldsymbol{1.46\cdot 10^{-1}}$ $\boldsymbol{4.05\cdot 10^{-3}}$ $\boldsymbol{4.18\cdot 10^{-3}}$ $4.17\cdot 10^{-1}$ $\boldsymbol{2.09\cdot 10^{-1}}$
Compressor LRU $1.88\cdot 10^{-2}$ $6.38\cdot 10^{-3}$ $5.49\cdot 10^{-3}$ $4.38\cdot 10^{-1}$ $4.52\cdot 10^{-1}$
S4D $2.00\cdot 10^{-1}$ $5.02\cdot 10^{-3}$ $6.12\cdot 10^{-3}$ $\boldsymbol{2.79\cdot 10^{-1}}$ $2.93\cdot 10^{-1}$
S6 $1.67\cdot 10^{-1}$ $4.41\cdot 10^{-3}$ $6.23\cdot 10^{-3}$ $4.99\cdot 10^{-1}$ $2.36\cdot 10^{-1}$
Neutron LSTM $1.06$ $3.15\cdot 10^{-2}$ $3.03\cdot 10^{-2}$ $9.33\cdot 10^{-1}$ $6.57\cdot 10^{-1}$
Filter ED $8.77\cdot 10^{-1}$ $2.18\cdot 10^{-2}$ $2.21\cdot 10^{-2}$ $9.38\cdot 10^{-1}$ $6.32\cdot 10^{-1}$
LRU $\boldsymbol{6.58\cdot 10^{-1}}$ $\boldsymbol{1.66\cdot 10^{-2}}$ $\boldsymbol{1.95\cdot 10^{-2}}$ $\boldsymbol{7.51\cdot 10^{-1}}$ $\boldsymbol{5.98\cdot 10^{-1}}$
S4D $7.18\cdot 10^{-1}$ $1.97\cdot 10^{-2}$ $2.00\cdot 10^{-2}$ $9.01\cdot 10^{-1}$ $6.28\cdot 10^{-1}$
S6 $8.23\cdot 10^{-1}$ $2.17\cdot 10^{-2}$ $2.18\cdot 10^{-2}$ $8.94\cdot 10^{-1}$ $6.27\cdot 10^{-1}$
In the OD300 and Neutron’s overdrive dataset, the LSTM model presents the lowest MSE and ESR, indicating better general amplitude matching between target and prediction. On the other hand, the S4D and S6 present better accuracy when considering the transients, the RMS envelope, and frequency-related content. This can be observed from the general lower error in terms of RMSE, spectral flux, STFT, and MFCC. In the case of maximum distortion, we can notice LRU prediction having significantly lower energy than other models. Notably, the spectrograms shown in Figure 5 illustrate that S4D and S6 models better match the energy in the harmonics at higher frequency regions. LRU fails to generate the harmonics in the OD300 case while ED presents the opposite behavior in Neutron’s overdrive. Models trained on the Neutron OD dataset exhibit larger errors, suggesting that this dataset presents a more challenging modeling case for the investigated architectures. This challenge can also be seen in the figures. The overdrive appears to be cutting the higher frequencies, behaving also as a low-pass filter.
For models trained on the Saturator dataset, the LRU model presents lower MSE and $M_{RMSE}$ values, but when considering other metrics, no models appear clearly outperforming. When inspecting Figures  4, LRU is again left behind, while other models present closer performances. In contrast, LRU achieves closer matching in the spectrograms, visible in Figure 5. Other models tend to generate more harmonic, especially LSTM that introduces artifacts, which can be clearly heard when listening to the audio output and negatively affect its performance.
When training models with the Equalizer dataset, LSTM performs better in the time domain, having lower MSE and ESR errors while performing poorly in the frequency domain. Figures  4 show close performance among the models, while looking at the spectrograms S6 present better matching. The optical compressor features a time-variant response with pronounced temporal dependencies, largely influenced by its variable attack and release parameters. This is especially noticeable in the case of the CL 1B compressor, which represents a more challenging task for neural network modeling. Among the tested models, the LSTM and LRU are the least accurate in the time domain metrics, exhibiting higher loss in both the LA-2A and CL 1B cases. In contrast, in the frequency domain, the gap is less prominent. The S6 and ED models demonstrate superior accuracy for these two tasks. Specifically, S6 outperforms the ED in the LA-2A case, while the ED excels in the CL 1B case. S4D presents similar performance than S6.
In the LA-2A case, we observe an overall good match for RMSE in Figure 4. The LRU model displays slightly more mismatches in both plots. Upon examining Figure  5, spectrograms related to the S6 model have the best match with the target output, while the other models present similar mismatches when looking at the higher frequencies. More difficult is the CL 1B case, where LRU presents the best envelope matching in Figure 4, and Figure  5. In the case of the compressor, the capability to encode past information of the signal is critical, LSTM is confirmed to have a worse temporal tracking ability than the other models, which present design improvements. These improvements are particularly evident with the CL 1B when setting long attack and release time through the conditioning control parameters. These reach up to $300$ ms and $5$ seconds, respectively.
Modeling the low-pass filter is inherently a more complex task than the equalizer, given its more pronounced alterations to the frequency content of the processed signal. Nevertheless, informal listening evaluations of the predicted outputs indicate that all models struggle with accurately emulating the filter behavior; instead of eliminating high-frequency contents, they attenuate the whole signal, as apparent in Figure 5. However, when considering the quantitative metrics, the LRU model outperforms the others. The LSTM model, in particular, tends to incorrectly predict excessive energy on the RMS envelope, as visible in Figure 4, and to introduce artifact, as shown in Figure 5.
Generally, performance metrics are relatively close across models across most effects examined in this study. Although the S4D model slightly outperforms other models based on the metrics alone, it appears to be the most suitable architecture analog effect modeling, at least considering those within the scope of this study. The S6 model often performs very closely, while LRU presents the bigger variance across the different metrics and datasets.
The ED shows its quality when modeling the compressors. The ED, LRU, and particularly the S4D and S6 models are more effective at encoding past signal information within their internal states. This is suggested by the case of the cases of optical compressors, where they significantly outperform other models. Conversely, the LSTM models appear to perform well only when emulating effects with shorter and less critical time dependencies, such as overdrive, equalizer, and filters. Moreover, LSTM models tend to introduce artifacts in some cases.
@@ 5 Conclusion
The computational challenges of simulating the nonlinearities of analog effect, described by complex differential equations, are a significant obstacle for real-time applications. The complexity of these nonlinearities makes them difficult to model accurately, but they also contribute to the unique sound characteristics of the devices, making them appealing for both academic and industrial purposes. Modeling these nonlinear aspects is a difficult task that requires simplifications and assumptions to ensure stability and efficiency in the numerical schemes. Artificial neural networks (ANN) have emerged as a popular data-driven approach for this modeling process. While they offer a convenient and automated approach, they may struggle with the computational complexity and flexibility required for handling control parameters. In some cases, hybrid solutions that incorporate physics knowledge or measurements from internal components may provide advantages. However, there are situations where black-box models are necessary and feasible. Recurrent neural networks (RNNs) utilize internal states as memory mechanisms, allowing them to encode the signal history. This feature enables more interactive solutions as their memory capacity is not solely dependent on the input amount. In this paper, we present a comparative analysis that examines the performance of the widely used Long Short-Term Memory (LSTM) models against recently proposed recurrent models, namely Linear Recurrent Unit (LRU) and State-Space models (SSM), which have demonstrated remarkable results in sequence modeling. To carry out this study, we design five models incorporating parameter conditioning and providing low-latency response. The control parameters are integrated into the models using the Feature-wise Linear Modulation (FiLM) method. Our selection of effects includes overdrive, saturation, low pass filter, equalization, and compressors. We evaluate these models using various metrics, which provide valuable insights for future research and development. Specifically, our metrics consider the root-mean-squared energy envelopes, the transients, and the frequency content.
LSTM-based networks excel in replicating distortion and equalizer effects, showing their ability to encode the audio dependencies enough for these specific effects effectively. Conversely, LSTMs tend to introduce more artifacts than other models considered in this study. In contrast, LSTM, when designed in an encoder-decoder fashion, LRUs and, in particular, SSMs, outperform other models when emulating saturation and compression. In the case of compression, they demonstrate their advantage in accurately capturing the signal’s history. Generally, SSMs have also proved to be an appropriate candidate technique in VA, although the gap with the other models is not always so pronounced. Lastly, LRUs are less consistent in their quality across the different metrics.